A charged paint is spread in a very thin uniform layer over the surface of a plastic sphere of diameter 20.0cm, giving it a charge of -10.0μC
A. Find the electric field just inside the paint layer.
B. Find the electric field just outside the paint layer.
C. Find the electric field 8.00cm outside the surface of the paint layer.
Answer #1

Given,

Diameter of the Sphere = 20 cm

Radius of the Sphere(R) = 10 cm = 0.1 m

Spherically symmetric charge distribution,Charge (q) = - 10.0μC = - 10 * 10-6 C = - 10-5 C

A. The electric field just inside the paint layer -(At an internal point)

Enclosed Charge = 0 implies, Electric Field = 0

B.The electric field just outside the paint layer-( At a point on the surface of sphere)

We know,

E = rac{1}{4pi epsilon _{0}} rac{q}{R^{2}}  

  -10 10-6 0.12 43,14x 8,85 x 10-12

E = 8.99 * 10) * (-10 * 10-4)

  E -8.99 * 10°ус

C.The electric field 8.00 cm outside the surface of the paint layer-(At an external point, r > R)

r = 8 cm + R = 8 cm + 10 cm = 18 cm = 0.18 m

E = rac{1}{4pi epsilon _{0}} rac{q}{r^{2}}

  一10 * 10-6 43,14 8,85 10-12 0.182

  E = 8.99 * 109 * (-30.86 * 10-5)

E-一277.43 * 101

E =- 2.77 * 10^{^{6}} N/C